OpenStudy (anonymous):

If sum of 4 consecutive natural number is 182 then find the numbers!

OpenStudy (anonymous):

44, 45, 46, 47

OpenStudy (anonymous):

thanks a looooot!

OpenStudy (turingtest):

four consecutive natural numbers means n+(n+1)+(n+2)+(n+3)=182 solve for n and see if Blank is right...

OpenStudy (turingtest):

I'm just demonstrating the concept. How did you solve it?

OpenStudy (anonymous):

just add four numbers to get 182

OpenStudy (anonymous):

yes,but i need a way to find them! anw,thanks

OpenStudy (anonymous):

i just guess and check

OpenStudy (anonymous):

I found that 44+45+46+47 =182

OpenStudy (anonymous):

yes,it's easy that way!

OpenStudy (turingtest):

ok, then smart guy, do that trick for 8 consecutive numbers that add to 7528.

OpenStudy (turingtest):

didn't intend to sound so mean, sorry. Just making a point.

OpenStudy (anonymous):

just another question! i dont know how to solve the way you told turingtest ? can u explain it please?

OpenStudy (turingtest):

are you asking me my method? or Blanks?

OpenStudy (anonymous):

yours!

OpenStudy (turingtest):

My method is: If you need to find x consecutive numbers that add up to y, then write\[n+(n+1)+(n+2)+...+(n+x-1)=y\]since you are given x and y you just need to solve for n. doing so will give you the value of n; the first number in the sequence. To find the other numbers just evaluate each set of parentheses. That will give you your sequence. Let's do it for your problem, where x=4 and y=182 n+(n+1)+(n+2)+(n+3)=182 4n+6=182 4n=176 n=44 so the numbers are 44, 44+1, 44+2, 44+3 =44,45,46,47

OpenStudy (turingtest):

get it?

OpenStudy (anonymous):

oooo thenks a lot! you're genius ahahh